A) A pulley has a string holding two masses which both hang straight down. If the mass on the left is 1.8 kg and the mass right is 3.5 kg, find the tension in the string.

B) A 5 kg block on a frictionless table is pulled to the right by a 15 N force. A string runs to the left to a second block that has a mass of 9. Find the acceleration.
Answer #1

A )

given

m1 = 1.8 kg

m2 = 3.5 kg

g = 9.8 m/sec2

T = ?

T = 2 g m1 m2 / ( m1 + m2 )

= 2 x 9.8 x 1.8 x 3.5 / ( 1.8 + 3.5 )

the tension in the string is T = 23.298 N

B )

using F = ma

a = F / ( M + m )

= 15 / ( 9 + 5 )

a = 15 / 14

the acceleration is a = 1.0714 m/sec2